A credit card company claims that the mean credit debt for individuals is about $5,000. To test this claim, a (type in left, right, or two) [a]-tailed test can be set up.

A random sample of 35 card holders has a mean balance $5,150 with a standard deviation $575. What is the P-value of this sample? Round to the nearest hundredth.

[b]

Is there enough evidence to reject the claim at the significant level 0.05?

[c] (type in yes or no)
Answer #1

Solution:

Two tailed test can be set up Test statistic, t = (xbar-mu)/(s/sqrt(n)) t = (5150 - 5000)/(575/sqrt(35)) t = 1.54 P-value App